1) When the demand function is 2Q - 24 + 3P = 0, calculate the price elasticity of demand when P = 4 and find the marginal revenue when Q=3.


2) A market consists of three people, A, B , and C, whose individual demand equation are as follows:

A: P= 35 - 0.5QA

B: P= 50 - 0.25QB

C: P= 40 - 2.00QC

The industry supply equation is given by Q= 40 + 3.5P

a) Determine the equilibrium price and quantity.

b) Determine the amount that will be purchased by each individual.


1
Expert's answer
2022-04-04T09:15:11-0400

1)

2Q-24+3P=0

Q=12-1.5P

"P=8-\frac{2}{3}Q"

"\epsilon=\frac{dQ}{dP}×\frac{P}{Q}"

Where "\frac{dQ}{dP}=-1.5"

"\epsilon=-1.5×\frac{4}{6}"

"\epsilon=-1"


"MR=\frac{dTR}{dQ}"

"P=8-\frac{2}{3}Q"

"TR=P\u00d7Q"

"TR=8Q-\frac{2}{3}Q^2"

"MR=8-\frac{4}{3}Q"


2)

P=35-0.5QA ; QA=70-2P

P=50-0.25QB ;. QB=200-4P

P=40-2QC ; QC=20-0.5P


Qs=40+3.5P

"E=Qd=Qs"

Where

"Qd=Q_A+Q_B+Q_C"

"Qd=290-6.5P"

"\therefore 290-6.5P=40+3.5P"

"250=10P"

"P^*=25"

"Q^*=290-6.5(25)=127.5"


"Q_A=70-2(25)"

"Q_A=20"


"Q_B=200-4(25)"

"Q_B=100"


"Q_C=20-0.5(25)"

"Q_C=7.5"